OpenStudy (anonymous):

bayside insurance offers two health plans. Under plan a giselle would have to pay the first 190 of her medical bills plus 35% of the rest. under plan b, giselle would pay the first $230, but only 25% of the rest. For what amount of medical bill will plan b save giselle money? Assume she has over $230 in bills

OpenStudy (anonymous):

this is not for the OS Feedback group

OpenStudy (anonymous):

Let's try and view this problem as 2 different sides of a divide. On one side, where plan A shines, we have the dollar values between $190 and $230, a total of $40 that plan B pays full price for and plan A only pays 35%, or $14. This saves $26 for Giselle. However, from here on out plan B takes over. For every dollar spent, plan A pays 35% and plan B pays 25%. So we need to know: .25x + 26 = .35x 26 = .10x 260 = x And don't forget this was 260 + where plan B takes over, the assumed $230. Everything over $490 with plan B saves Giselle money.

OpenStudy (amistre64):

G pays = constant + (bill-constant)% = constant + bill(%) - constant(%) = bill% + constant(1) - constant(%) = bill(%) + constant(1-%) planA: = Bill(.35) +190(.65) = Bill(.35) +123.50 planB = Bill(.25) +230(.75) = Bill(.25) +172.50 Bill(.35) +123.50 = Bill(.25) +172.50 -Bill(.25) - 123.50 -Bill(.25) - 123.50 ------------------------------------- Bill(.10) = 49 Bill = 49/.1 = $490